OpenStudy (anonymous):

A right triangle has one leg with a length of 48 and a hypotenuse with a length of 80. What is the length of the other leg?

OpenStudy (anonymous):

We can either use the Pythagorean theorem to solve this problem, or we can use trig functions. To solve using the Pythagorean theorem, plug in values into the formula \[a^2 + b^2 = c^2\] the formula with the values plugged in would be \[48^2 + b^2 = 80^2\] which would then be \[2,304 + b^2 = 6,400\] do you think you can solve the equation now?

OpenStudy (anonymous):

If you don't know how to solve after that, subtract 2,304 from both sides to get \[b^2 = 4,096\] then square root both sides to get \[b = \sqrt{4096}\] I'll leave the rest to you

OpenStudy (midhun.madhu1987):